# Differential Privacy of Cross-Attention with Provable Guarantee

### v2

@@ B.4 Comparison of Truncated Laplace, Gaussian, and Laplace Mechanisms
Definition B.9 (Laplace distribution).
@@
In this paper, we use the Chebyshev inequality to bound the error, and from GDGK [32], we know that the truncated Laplace mechanism has the current minimum variance across all $(\epsilon,\delta)$-DP distributions.
The variance of Gaussian mechanism in Theorem 3.22 in DR [20]:
@@
The variance of Laplace mechanism in Fact B.10:
@@
The variance of truncated Laplace mechanism in Fact 2.5, for $c\in(0,1]$:
@@
Appendix C Proof Outline
This section provides the proof outline of our paper. In Section C.1, we analyze our DPTree data structure. In Section C.2, we show the sensitivity of summation problem. In Section C.3, we explain the high-level idea behind the weighted $\ell_{p}^{p}$ distance query. In Section C.4, we show how to answer one-dimensional weighted $\ell_{1}$ distance query. In Section C.5, we show how to answer Softmax distance query using previous algorithms. In Section C.6, we show how to handle adaptive query. By combining the results from these sections, we prove the main results in Section 4.
C.1 Summation Segment Tree
First, in order to solve the weighted distance problem, we need to have a basic DP algorithm (Algorithm 2) that can answer simple summation queries. After analyzing its DP and error in Section D, we state the data structure theorem.
Theorem C.1 (DPTree data structure, informal version of Theorem D.1).
@@ C.1 Summation Segment Tree
  • Init$(a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{N}_{+},\epsilon\in(0,1),\delta\in(0,1))$. It takes $O(n)$ time to initialize the data structure.
  • Query$(x\in[n],y\in[n])$. It takes $O(\log n)$ time to output a number $z$ such that
    • the process of output $z$ satisfies $(\epsilon,\delta)$-DP private, which computes $\sum_{i=x}^{y}a_{i}$,
    • $|z-\sum_{i=x}^{y}a_{i}|\leq O(\epsilon^{-1}\Delta\log^{3/2}n)$,
    • it holds with probability $0.99$.
During the design of the data structure, we found an interesting property based on the parallel composition rule of DP Fact B.7. We will now state the lemma, whose proof is provided in Section D.
Lemma C.2 (Weighted sum of disjoint interval queries, informal version of Lemma D.8).
If the following conditions hold that:
  • Let there be $t$ disjoint intervals, i.e., $S_{j}$ for $j\in[t]$, such that $S_{j}\cap S_{k}=\emptyset$ for all $j\neq k$.
  • Let $\epsilon\in(0,1)$ and $\delta\in(0,1)$.
  • Let $a_{j}$ for $j\in[t]$ be a series that square converges to $a$, i.e., $\sum_{j=1}^{t}a_{j}^{2}\leq a$.
Then, we have Alg. 2 is $(\epsilon,\delta)$-DP and output $\sum_{j=1}^{t}a_{j}\textsc{Query}(S_{j})$ with the error upper bounded by
$O(a^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)$
with probability $0.99$.
From Lemma C.2, we can see that if we have multiple disjoint interval queries, the error of the weighted sum of the intervals can be bounded independently of the number of queries, as long as the sum of squared weights is finite.
C.2 Sensitivity for Range Summation Problem
Lemma C.3 (Sensitivity of summation).
Proof.
where the first step follows from Definition 2.2, the second step follows from $X,X^{\prime}$ differ in $x_{i},x^{\prime}_{i}$, and the last step follows from each coordinate of the dataset is bounded in $[-R,R]$. ∎
C.3 Weighted $\ell_{p}^{p}$ Distance Problem
In this section, we introduce the intuition behind the method for handling the weighted $\ell_{p}^{p}$ distance problem. The formal lemmas and proofs can be found in Section E.
Given a dataset and a query point in $d$ dimensions, we round each coordinate of the data points and the query point to the nearest multiple of a small interval. We then aggregate the weights of data points that have been rounded to the same position. Finally, we compute the sum of these aggregated weights multiplied by the distances between the query point and the data points over the rounded positions. This approach makes the computation more efficient while maintaining sufficient accuracy.
We provide an example of weighted $\ell_{1}$-distance of a one-dimensional dataset consisting of 10 data points, i.e., $X\in[0,1]^{10}$ and a query $y=0$ in Figure 1.
Lemma C.4 (Weighted $\ell_{p}^{p}$-distance high dimension, informal version of Lemma E.2).
If the following conditions hold:
  • Let data $X\in[0,R]^{n\times d}$ and $x_{i}^{\top}\in[0,R]^{d}$ be the $i$-th row of $x$, weight $w\in\mathbb{R}^{n}$, query $y\in[0,R]^{d}$.
  • We round each dimension of $X$ and $y$ to an integer multiple of $R/n$.
  • Let $x_{i,k},y_{k}$ denote the $k$-th coordinates of $x_{i},y$ for $k\in[d]$.
  • Let $c_{j,k}:=\sum_{j_{0}\in S_{j,k}}w_{j_{0}}$ where the set $S_{j,k}$ is the set of index $i$ such that the corresponding $x_{i,k}$ is rounded to $jR/n$ for $j\in\{0,1,2,\dots,n\}$ for $k\in[d]$.
  • After rounding, we assume that $y_{k}$ is in the $l_{k}R/n$ position for $l_{k}\in\{0,1,2,\dots,n\}$ for $k\in[d]$.
For the weighted problem, we have
$\sum_{i=1}^{n}w_{i}\cdot\|y-x_{i}\|_{p}^{p}=\sum_{k=1}^{d}\sum_{j=0}^{n}(|l_{k}-j|R/n+O(R/n))^{p}c_{j,k}.$
where $O(R/n)$ is the rounding error for each data point.
Remark C.5.
In Lemma C.4, we first round the dataset. This rounding simplifies the calculation by reducing the number of possible positions to consider, from real values in $[0,R]^{d}$ to the total $O(nd)$ spots. However, it also introduces an error $O(R/n)$ for one data point. Then, for one spot in the rounded dataset, we sum over the weights of that spot and multiply the corresponding distance raised to the power of $p$. Additionally, since we are dealing with $\ell_{p}^{p}$ distance, the rounding error is also raised to the power of $p$.
Figure 1: The visualization of how to build $c_{j}$ of rounded dataset $X\in[0,1]^{10}$ and compute the weighted $\ell_{1}$ distance. The number above each $x_{i}$ is $w_{i}$. See Algorithm 5 for details. Suppose $y=0$. Then $\sum_{i=1}^{n}w_{i}|y-x_{i}|=0.1\cdot 2.2+0.3\cdot 3.1+0.3\cdot(-2)+0.3\cdot(-3)+0.4\cdot 2+0.6\cdot 6+0.7\cdot 0.5+0.9\cdot(-1)+0.9\cdot 1=4.4$. And $\sum_{j=0}^{n}|k-j|c_{j}/n=0.1\cdot 2.2+0.3\cdot(-1.9)+0.4\cdot 2+0.6\cdot 6+0.7\cdot 0.5=4.4$. See details in Lemma E.1.
C.4 One-Dimensional Weighted $\ell_{1}$ Distance Data Structure
Based on previous discussions in Section C.1 and C.3, we can now describe our one-dimensional weighted $\ell_{1}$ distance data structure, DPTreeDistance, presented in Algorithm 5 and 6, which generalizes the results from BLM+ [10]. Drawing from the intuition in Section C.3, the initialization process is as follows: first, we round each data point in the dataset to the nearest multiple of a small interval and build an array that aggregates the corresponding weights. This array is then fed into our DPTree data structure in Algorithm 2. At query time, we query the DPTree to obtain the aggregated weights within a small interval and multiply these weights by the distance to the query point. Furthermore, we also introduce a relative error parameter $\alpha$ to reduce the total number of queries to $O(\log(n)/\alpha)$ instead of querying all $n$ positions. We also analyze the DP and the error bound; see details in Section F.
Theorem C.6 (DPTreeDistance data structure, informal version of Theorem F.6).
There is a data structure DPTreeDistance (Algorithm 5 and 6) that uses $O(n)$ spaces to solve weighted $\ell_{1}$-distance query problem for dataset $X\subset[0,R]$ and support the following operations:
  • Init$(X\subset[0,R],n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1))$. (Algorithm 5) It takes $O(n)$ time to initialize the data structure.
  • DistanceQuery$(y\in[0,R],\alpha\in(0,1))$. (Algorithm 6) It takes $O(\alpha^{-1}\log^{2}n)$ time to output a number $z$ such that
    • the process of output $z$ satisfies $(\epsilon,\delta)$-DP, which computes $\sum_{i\in[n]}w_{i}|y-x_{i}|$,
    • $|z-\sum_{i\in[n]}w_{i}|y-x_{i}||\leq\alpha\sum_{i\in[n]}w_{i}|y-x_{i}|+O(\epsilon^{-1}\alpha^{-1/2}RR_{w}\log^{3/2}n)$,
    • it holds with probability $0.99$.
C.5 Softmax Activation
We then describe how we extend the previous results to Softmax activation, i.e. exponential inner product function (Definition 4.1). From AS [6], we know that Softmax activation can be approximated by polynomial kernel function $P(\cdot)$ with a certain error. The following lemma shows that we can transform weighted Softmax queries into polynomial kernels. More specifically, we have one term that computes the weighted $\ell_{2}^{2}$ distance, which is the place where we add DP noises. Because of the decomposability of the $\ell_{p}^{p}$ distance, i.e. $\sum_{i\in[n]}w_{i}\|x_{i}-y\|_{p}^{p}=\sum_{j\in[d]}\sum_{i\in[n]}w_{i}|x_{i,j}-y_{j}|^{p}$, we can easily extend the results of Section C.4 to handle the $\ell_{2}^{2}$ distance query. After that, we compute the term for the weighted $\ell_{2}^{2}$ norms of approximation kernel exactly. Summing all these terms, with a certain error, we can answer the Softmax query. Related proofs can be found in Section I.
Lemma C.7 (Weighted Softmax approximation, informal version of Lemma I.6).
Let the accuracy parameter be $\epsilon_{s}\in(0,0.1)$. Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition I.3). Let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in Lemma I.5. Then, we can approximate the exponential inner product using the polynomial kernel function:
$w^{\top}\exp(Xy/d)=$ $~{}-\frac{1}{2}\sum_{j\in[r]}\sum_{i\in[n]}w_{i}|P(x_{i})_{j}-P(y)_{j}|^{2}+\frac{1}{2}\sum_{i\in[n]}w_{i}(\|P(x_{i})\|^{2}_{2}+\|P(y)\|^{2}_{2})$
$+~{}O(w^{\top}\exp(Xy/d)\cdot\epsilon_{s}).$
Moreover, the vectors $P(\cdot)$ can be computed in $O(r)$ time.
C.6 Adaptive Query
We introduce how we can modify our algorithm to solve the adaptive query problem using some tools in QRS+ [86]. Our approach is based on proving that our algorithm can handle any query within the query space with a certain error. Since adaptive queries must lie within this space, our algorithm can effectively handle them. In Section C.5, we demonstrate our algorithm’s capability to answer weighted Softmax distance queries with constant probability. We then use the Chernoff bound to boost the constant probability of our algorithm to a high probability. Next, we apply the notion of an $\epsilon_{0}$-net to bound all query points within the net using the union bound. Finally, we bound all points in the query space by utilizing the Lipschitz property of the weighted Softmax distance function and introducing an additive error. See the proofs in Sections H and I.
Lemma C.8 (Adaptive Softmax, informal version of Lemma I.10).
If the following conditions hold:
  • Let $N$ be the $\ell_{\infty}$ $\epsilon_{0}$-net of ${\mathcal{B}}$, and let $|N|$ be the size of the net $N$.
  • Let the relative error parameter $\alpha\in(0,1)$, the failure probability $p_{f}\in(0,0.01)$.
  • We create $l=O(\log((R/\epsilon_{0})^{r}/p_{f}))$ independent copies of the data structure $\{\textsc{DPTreeSoftmax}_{j}\}_{j=1}^{l}$ (Algorithm 3) and take the median of the outputs with each data structure instantiated with $(\epsilon/l,(\delta+\delta^{\prime})/l)$-DP.
  • Let $f(y):=\mathrm{Median}(\{\textsc{DPTreeSoftmax}_{j}.\textsc{DistanceQuery}(y,\alpha)\}_{j=1}^{l})$.
  • Let $Z(y):=w^{\top}\exp(Xy/d)$.
  • Let $B=O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
Then with probability $1-p_{f}$, for all query points $q\in{\mathcal{B}}$, there exists a point $y\in N$ which is the closest to $q$, we can have the process of outputting median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error satisfies
$|f(y)-Z(q)|\leq(\alpha+\epsilon_{s})Z(q)+B+2n\sqrt{d}RR_{w}\exp(R^{2})\epsilon_{0}.$
Appendix D DPTree Algorithm
In this section, we give the analysis of privacy, accuracy and runtime of our DPTree (Algorithm 2). In Section D.1, we give the theorem (Theorem D.1) of our data structure that can answer summation problem. In Section D.2, we improve our data structure from constant probability to high probability by applying Chernoff bound. In Section D.3, we give the analysis. In Section D.4, we show some results of our data structure if the input queries are disjoint.
D.1 Single Data Structure
Theorem D.1 (DPTree data structure, formal version of Theorem C.1).
  • Init$(a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{N}_{+},\epsilon\in(0,1),\delta\in(0,1))$. It takes $O(n)$ time to initialize the data structure.
  • Query$(x\in[n],y\in[n])$. It takes $O(\log n)$ time to output a number $z$ such that
    • the process of output $z$ satisfies $(\epsilon,\delta)$-DP private, which computes $\sum_{i=x}^{y}a_{i}$,
    • $|z-\sum_{i=x}^{y}a_{i}|\leq O(\epsilon^{-1}\Delta\log^{3/2}n)$,
@@ D.1 Single Data Structure
The proofs follow from combining Lemma D.3 (running time of initialization), Lemma D.4 (running time of query), Lemma D.5 (DP of query), and Lemma D.6 (error of query) together. ∎
D.2 Boost the Constant Probability to High Probability
We can use Chernoff bound to boost the high probability by repeating the data structure multiple times.
Theorem D.2 (High-probability).
There is a data structure (see DPTreeHighProb in Algorithm 4) that uses $O(n\log(1/\delta_{\mathrm{fail}}))$ spaces to support the following operations
  • Init$(a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{N}_{+},\epsilon\in(0,1),\delta\in(0,1),\delta_{\mathrm{fail}}\in(0,0.01))$. It takes $O(n\log(1/\delta_{\mathrm{fail}}))$ time to initialize the data structure.
  • Query$(x\in[n],y\in[n])$. It takes $O(\log(n)\cdot\log(1/\delta_{\mathrm{fail}}))$ time to output a number $z$ such that
    • the process of output $z$ satisfies $(\epsilon,\delta)$-DP private, which computes $\sum_{i=x}^{y}a_{i}$,
    • $|z-\sum_{i=x}^{y}a_{i}|\leq O(\epsilon^{-1}\Delta\log^{3/2}(n)\cdot\log(1/\delta_{\mathrm{fail}}))$,
@@ D.2 Boost the Constant Probability to High Probability
Note that our data structure (Theorem D.1) succeeds with probability $0.99$. The success of the algorithm (Theorem D.1) can be viewed as a Bernoulli random variable, to which we apply the Chernoff bound (Lemma B.2). By repeating the data structure $O(\log(1/\delta_{\mathrm{fail}}))$ times and taking the median of the outputs, we boost the success probability. The details are following.
To boost the success probability, we assume the query is repeated $l$ times. Let $i\in[l]$, and let $z_{i}$ denote the indicator random variable for the success of the $i$-th instance of the data structure for a single query. Let $z=\sum_{i=1}^{l}z_{i}$ be the total success times. Since $p=\Pr[z_{i}=1]=0.99$, we can have $\mu=\operatorname*{{\mathbb{E}}}[z]=\sum_{i=1}^{l}p=lp$. Note that $p=0.99$. By setting $\delta=0.1$ and using Chernoff bound from Lemma B.2, we can show
@@
By the DP basic composition rule (Fact B.6), we need to choose $\epsilon=\epsilon^{\prime}/O(\log(1/\delta_{\mathrm{fail}}))$ and $\delta=\delta^{\prime}/O(\log(1/\delta_{\mathrm{fail}}))$ where $\epsilon^{\prime},\delta^{\prime}$ are the $\epsilon,\delta$ in Theorem D.1. ∎
Algorithm 4 Boost constant probability
1:datastructure DPTreeHighProb $\triangleright$ Theorem D.2
2:members
3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{O(\log(1/\delta_{\mathrm{fail}}))}:\textsc{DPTree}$triangleright Alg. 2
4:end members
5:procedure Init($a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{N}_{+},\epsilon\in(0,1),\delta\in(0,1),\delta_{\mathrm{fail}}\in(0,0.01)$)
6:     for $i=1\to O(\log(1/\delta_{\mathrm{fail}}))$ do
7:         ${\mathcal{D}}_{i}$.Init$(a,n,\Delta,\epsilon/O(\log(1/\delta_{\mathrm{fail}})),\delta/O(\log(1/\delta_{\mathrm{fail}})))$
8:     end for
9:end procedure
10:procedure Query($x\in[n],y\in[n]$)
11:     $r\leftarrow 0^{O(\log(1/\delta_{\mathrm{fail}}))}$
12:     for $i=1\to O(\log(1/\delta_{\mathrm{fail}}))$ do
13:         $r_{i}\leftarrow{\mathcal{D}}_{i}.\textsc{Query}(x,y)$
14:     end for
15:     return Median of $r$
16:end procedure
17:end datastructure
D.3 Algorithm of Data Structure
@@
Lemma D.3 (Runtime of initialization, Algorithm 2).
@@
Lemma D.4 (Runtime of query, Algorithm 2).
@@
Lemma D.5 (Privacy of query, Algorithm 2).
@@
Suppose that our dataset is $X\in[-R,R]^{n}$. Note that we only add noise in the pre-processing stage. There is no noise in the query stage. Since the problem we care about is summation, if we change one leaf node, the sensitivity $\Delta=2R$ (see Lemma C.3). Since we add noise to each node in the tree, and each leaf node count will contribute to $\log n$ nodes, it is equivalent to our output function being in $\log n$ dimension. We will then blow up the DP parameter by $\log n$ factor. Thus, using the basic composition rule (Fact B.6), the DP guarantee for the whole tree data structure is $((\epsilon/\log n)\cdot\log n,(\delta/\log n)\cdot\log n)$ which is $(\epsilon,\delta)$-DP. ∎
@@
Lemma D.6 (Accuracy of query, Algorithm 2).
Let $\epsilon\in(0,1)$ and $\delta\in(0,1)$. Then, using Chebyshev’s inequality and Fact 2.5, we have the error of Query(see Algorithm 2) output is upper bounded by:
@@
[…]
@@
We know $\mu=\operatorname*{{\mathbb{E}}}[z]=0$ since $\operatorname*{{\mathbb{E}}}[z_{i}]=0$. From Fact 2.5, we know the variance for each $z_{i}$ is $\operatorname*{\mathrm{Var}}[z_{i}]=c\epsilon^{-2}\Delta^{2}\log^{2}n$ where $0<c\leq 2$ and $c=2$ when $\delta=0$.
@@
Note that we wish to bound $|z|=|\textsc{Query}(S)-\textsc{TrueQuery}(S)|$ as our error.
Using Lemma B.3, we can have
@@
D.4 Disjoint Intervals
In this section, we show some interesting results for our DPTree data structure if the input queries are disjoint.
Lemma D.7 (Sum of disjoint interval queries).
If the following conditions hold that:
  • Let there be $t$ disjoint intervals, i.e., $S_{j}$ for $j\in[t]$, such that $S_{j}\cap S_{k}=\emptyset$ for all $j\neq k$.
  • Let $\epsilon\in(0,1)$ and $\delta\in(0,1)$.
Then, we have Algorithm 2 is $(\epsilon,\delta)$-DP and outputs $\sum_{j=1}^{t}\textsc{Query}(S_{j})$ with the error upper bounded by
$O(t^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)$
with probability $0.99$.
Proof.
From Lemma D.5, we know that $\mathrm{DPTree}$.Query is $(\epsilon,\delta)$-DP. Then, from Fact B.7 and the disjoint intervals in Algorithm 6, we can conclude that the value returned is $(\epsilon,\delta)$-DP.
Let $\textsc{TrueQuery}(S_{j})$ denote the true interval query answer returned by DPTree.TrueQuery in Algorithm 2 for interval $S_{j}$. Let $\textsc{Query}(S_{j})$ denote the noised interval query answer returned by DPTree.Query in Algorithm 2 for interval $S_{j}$. Let $z_{j}:=\textsc{Query}(S_{j})-\textsc{TrueQuery}(S_{j})$ and $z=\sum_{j=1}^{t}z_{j}$. From the proof of Lemma D.6, we know $z_{j}$ is the sum of $O(\log n)$ independent truncated Laplace random variables each with parameter $\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$ and the variance is bounded by
$\operatorname*{\mathrm{Var}}[z_{j}]=O(\epsilon^{-2}\Delta^{2}\log^{3}n)$
Since the intervals $S_{j}$ are disjoint, they are independent to each other. Then, we have
$\operatorname*{\mathrm{Var}}[z]=$ $~{}\operatorname*{\mathrm{Var}}[\sum_{j=1}^{t}z_{j}]$
$=$ $~{}\sum_{j=1}^{t}\operatorname*{\mathrm{Var}}[z_{j}]$
$=$ $~{}O(t\epsilon^{-2}\Delta^{2}\log^{3}n)$
where the first step follows from definition of $z$, the second step follows from the intervals are disjoint, and the last step follows from $\operatorname*{\mathrm{Var}}[z_{j}]=O(\epsilon^{-2}\Delta^{2}\log^{2}n)$.
Note that we wish to bound $|z|$ as our error.
Using Lemma B.3, we can have error bounded by
$O(t^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)$
with probability $0.99$. ∎
Moreover, this can be generalized to weighted sum of queries.
Lemma D.8 (Weighted sum of disjoint interval queries, formal version of Lemma C.2).
If the following conditions hold that:
  • Let there be $t$ disjoint intervals, i.e., $S_{j}$ for $j\in[t]$, such that $S_{j}\cap S_{k}=\emptyset$ for all $j\neq k$.
  • Let $\epsilon\in(0,1)$ and $\delta\in(0,1)$.
  • Let $a_{j}$ for $j\in[t]$ be a series that square converges to $a$, i.e., $\sum_{j=1}^{t}a_{j}^{2}\leq a$.
Then, we have Alg. 2 is $(\epsilon,\delta)$-DP and output $\sum_{j=1}^{t}a_{j}\textsc{Query}(S_{j})$ with the error upper bounded by
$O(a^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)$
with probability $0.99$.
Proof.
The DP proof is the same as in the proof of Lemma D.7.
Let $\textsc{TrueQuery}(S_{j})$ and $\textsc{Query}(S_{j})$ be same in the proof of Lemma D.7 Let $z_{j}:=\textsc{Query}(S_{j})-\textsc{TrueQuery}(S_{j})$ and $z=\sum_{j=1}^{t}a_{j}z_{j}$. From the proof of Lemma D.7, we know the variance of $z_{j}$ is bounded by
$\operatorname*{\mathrm{Var}}[z_{j}]=O(\epsilon^{-2}\Delta^{2}\log^{3}n)$
Since the intervals $S_{j}$ are disjoint, they are independent to each other. Then, we have
$\operatorname*{\mathrm{Var}}[z]=$ $~{}\operatorname*{\mathrm{Var}}[\sum_{j=1}^{t}a_{j}z_{j}]$
$=$ $~{}\sum_{j=1}^{t}\operatorname*{\mathrm{Var}}[a_{j}z_{j}]$
$=$ $~{}\sum_{j=1}^{t}a_{j}^{2}\operatorname*{\mathrm{Var}}[z_{j}]$
$=$ $~{}\sum_{j=1}^{t}a_{j}^{2}\cdot O(\epsilon^{-2}\Delta^{2}\log^{3}n)$
$=$ $~{}O(a\epsilon^{-2}\Delta^{2}\log^{3}n)$
where the first step follows from the definition of $z$, the second step follows from the intervals are disjoint, the third step follows from the $\operatorname*{\mathrm{Var}}[az]=a^{2}\operatorname*{\mathrm{Var}}[z]$ for a random variable $z$ and a constant $a$, the fourth step follows from the $\operatorname*{\mathrm{Var}}[z_{j}]=O(\epsilon^{-2}\Delta^{2}\log^{2}n)$, and the last step follows from $\sum_{j=1}^{t}a_{j}^{2}\leq a$.
Note that we wish to bound $|z|$ as our error.
Using Lemma B.3, we can have error bounded by
$O(a^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)$
with probability $0.99$. ∎
Appendix E Weighted $\ell_{p}^{p}$ Distance
In this section, we introduce how to handle weighted $\ell_{p}^{p}$ distance problem in the high level idea. In Section E.1, we show how to solve one dimensional weighted problem. In Section E.2, we show how to solve high dimensional weighted problem by decomposing each coordinate of the high dimensional dataset.
Suppose we have the original data $X\in[0,R]^{n}$ and weight $w\in\mathbb{R}^{n}$ and query $y\in[0,R]$. We want to compute the weighted $\ell_{1}$-distance, i.e.
$\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|.$

### v3

@@ A.4 Comparison of Truncated Laplace, Gaussian, and Laplace Mechanisms
Definition A.11 (Laplace distribution).
@@
In this paper, we use the Chebyshev inequality to bound the error, and from [26], we know that the truncated Laplace mechanism has the current minimum variance across all $(\epsilon,\delta)$-DP distributions.
The variance of Gaussian mechanism in Theorem 3.22 in [17]:
@@
The variance of Laplace mechanism in Fact A.12:
@@
The variance of truncated Laplace mechanism in Fact 3.5, for $c\in(0,1]$:
@@
Appendix B DPTree Algorithm
In this section, we give the analysis of privacy, accuracy and runtime of our DPTree (Algorithm 2).
B.1 Single Data Structure
Theorem B.1 (DPTree data structure ).
@@ B.1 Single Data Structure
[…]
    • the process satisfies $(\epsilon,\delta)$-DP,
    • $|z_{1}-\sum_{\{k|x_{k}\leq y\}}a_{k}|\leq O(\epsilon^{-1}\Delta\log^{3/2}n)$ and $|z_{2}-\sum_{\{k|x_{k}\geq y\}}a_{k}|\leq O(\epsilon^{-1}\Delta\log^{3/2}n)$,
@@
The proofs follow from combining Lemma B.4 (running time of initialization), Lemma B.5 (running time of query), Lemma B.6 (DP of query), and Lemma B.7 (error of query) together. ∎
Figure 1: The visualization of how to compute the weighted $\ell_{1}$ distance for rounded dataset $X\in[0,1]^{10}$. The number above each $x_{i}$ is $w_{i}$. See Algorithm 4 for details. Suppose $y=0$. Then $\sum_{i=1}^{n}w_{i}|y-x_{i}|=0.1\cdot 2.2+0.3\cdot 3.1+0.3\cdot(-2)+0.3\cdot(-3)+0.4\cdot 2+0.6\cdot 6+0.7\cdot 0.5+0.9\cdot(-1)+0.9\cdot 1=4.4$. See more details in Lemma C.1.
B.2 Boost the Constant Probability to High Probability
By applying the Chernoff bound, we can increase the probability of obtaining a correct result. This is achieved by replicating the data structure multiple times, generating several independent results, and then reporting the median of these results. Taking the median helps mitigate the effect of outliers and ensures that the final answer is reliable with high probability.
Theorem B.2 (High-probability).
There is a data structure that uses $O(n\log(1/\delta_{\mathrm{fail}}))$ spaces to support the following operations
[…]
    • the process satisfies $(\epsilon,\delta)$-DP,
    • $|z_{1}-\sum_{\{k|x_{k}\leq y\}}a_{k}|\leq O(\epsilon^{-1}\Delta\log^{3/2}(n)\cdot\log(1/\delta_{\mathrm{fail}}))$ and $|z_{2}-\sum_{\{k|x_{k}\geq y\}}a_{k}|\leq O(\epsilon^{-1}\Delta\log^{3/2}(n)\cdot\log(1/\delta_{\mathrm{fail}}))$,
@@ B.2 Boost the Constant Probability to High Probability
Note that our data structure (Theorem B.1) succeeds with probability $0.99$. The success of the algorithm (Theorem B.1) can be viewed as a Bernoulli random variable, to which we apply the Chernoff bound (Lemma A.2). By repeating the data structure $O(\log(1/\delta_{\mathrm{fail}}))$ times and taking the median of the outputs, we boost the success probability. The details are following.
To boost the success probability, we assume the query is repeated $l$ times. Let $i\in[l]$, and let $z_{i}$ denote the indicator random variable for the success of the $i$-th instance of the data structure for a single query. Let $z=\sum_{i=1}^{l}z_{i}$ be the total success times. Since $p=\Pr[z_{i}=1]=0.99$, we can have $\mu=\operatorname*{{\mathbb{E}}}[z]=\sum_{i=1}^{l}p=lp$. Note that $p=0.99$. By setting $\delta=0.1$ and using Chernoff bound from Lemma A.2, we can show
@@
By the DP basic composition rule (Fact A.8), we need to choose $\epsilon=\epsilon^{\prime}/O(\log(1/\delta_{\mathrm{fail}}))$ and $\delta=\delta^{\prime}/O(\log(1/\delta_{\mathrm{fail}}))$ where $\epsilon^{\prime},\delta^{\prime}$ are the $\epsilon,\delta$ in Theorem B.1. ∎
B.3 Sensitivity for Summation Problem
Lemma B.3 (Sensitivity of summation).
Proof.
where the first step follows from Definition A.6, the second step follows from $X,X^{\prime}$ differ in $x_{i},x^{\prime}_{i}$, and the last step follows from each coordinate of the dataset is bounded in [R,R][-R,R]. ∎
B.4 Algorithm of Data Structure
@@
Lemma B.4 (Runtime of initialization, Algorithm 2).
@@
Lemma B.5 (Runtime of query, Algorithm 2).
@@
Lemma B.6 (Privacy of query, Algorithm 2).
@@
Suppose that our dataset is $X\in[-R,R]^{n}$. Note that we only add noise in the pre-processing stage. There is no noise in the query stage. Since the problem we care about is summation, if we change one leaf node, the sensitivity $\Delta=2R$ (see Lemma B.3). Since we add noise to each node in the tree, and each leaf node count will contribute to $\log n$ nodes, it is equivalent to our output function being in $\log n$ dimension. We will then blow up the DP parameter by $\log n$ factor. Thus, using the basic composition rule (Fact A.8), the DP guarantee for the whole tree data structure is $((\epsilon/\log n)\cdot\log n,(\delta/\log n)\cdot\log n)$ which is $(\epsilon,\delta)$-DP. ∎
@@
Lemma B.7 (Accuracy of query, Algorithm 2).
Let $\epsilon\in(0,1)$ and $\delta\in(0,1)$. Then, using Chebyshev’s inequality and Fact 3.5, we have the error of Query(see Algorithm 2) output is upper bounded by:
@@
Let $y\in[0,R]$ be the query.
Let $A_{1},A_{2}=\textsc{Query}(y)$ denote the noised query answers returned by DPTree.Query in Algorithm 2. Let $A^{*}_{1},A^{*}_{2}$ be the true query answers without noise. Let $z:=A_{1}-A_{1}^{*}+A_{2}-A_{2}^{*}$, which from Algorithm 2 we can see this is the sum of $O(\log n)$ independent truncated Laplace random variables each with parameter $\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$. Thus,
@@
We know $\mu=\operatorname*{{\mathbb{E}}}[z]=0$ since $\operatorname*{{\mathbb{E}}}[z_{i}]=0$. From Fact 3.5, we know the variance for each $z_{i}$ is $\operatorname*{\mathrm{Var}}[z_{i}]=c\epsilon^{-2}\Delta^{2}\log^{2}n$ where $0<c\leq 2$ and $c=2$ when $\delta=0$.
@@
Note that we wish to bound $|z|$ as our error.
Using Lemma A.3, we can have
@@
Appendix C Weighted $\ell_{p}^{p}$ Distance
In this section, we introduce how to handle weighted $\ell_{p}^{p}$ distance problem in the high level idea. We can solve high dimensional weighted problem by decomposing each coordinate of the high dimensional dataset. Thus, we only need to show how to solve the one-dimensional weighted problem.